Basil at $6.20 per ounce and oregano at $3.70 per ounce are mixed to obtain 20 ounces of blend worth $5.20 per ounce. How many ounces of each herb are used?

Found 2 solutions by dfrazzetto, josgarithmetic:
Answer by dfrazzetto(283)

Basil - 6.20 per ounce
Set ounces of basil = x
Total value of basil = 6.2x

Oregano - 3.70 per ounce
Set ounces of Oregano = y
Total value of Oregano = 3.7y
x + y = 20 ounces total
6.2x + 3.7y = 5.2(20)
x + y = 20; x = 20 - y
6.2(20 - y) + 3.7y = 104
124 - 6.2y + 3.7y = 104
124 - 2.5y = 104
-2.5y = -20
y = 8 ounces of Oregano
x = 20 - y = 12 ounces of Basil
CHECK:
8 + 12 = 20 ounces total
8(3.7) + 12(6.2) = 20(5.2)
104 = 104; Correct

Answer by josgarithmetic(39617)

You want this:
http://www.algebra.com/my/mixture-price-two-part-both-parts-unknown.lesson?content_action=show_dev